# Extreme events at the onset of epileptic-like chimeras in small-world networks of FitzHugh-Nagumo neurons

### v1

@@ Appendix A Numerical simulations
The dynamics of coupled FHN oscillators are solved by integrating Eq. (1) with a fourth-order Runge-Kutta method, considering random initial conditions following a uniform distribution $x\sim U_{x}(-a,a)$, $y\sim U_{y}(-a+a^{3}/3,a+a^{3}/3)$ and a suitable time step that ensures convergence. Specifically, simulations in (Fig. 2) consider the time evolution of $1\times 10^{4}$ steps, and the ISI PDF is calculated after a transient of $1\times 10^{2}$ time steps. Simulations in (Fig. 3a-e, 4a-d and 7a) consider the time evolution of $1\times 10^{5}$ time steps. The computation of $\Omega$ for each point in the $N-d-p$ space in (Fig. 3a-c) is calculated by averaging the analysis of 50 realisations of $R(t)$, excluding the final $1\times 10^{3}$ time steps, in order to consider only epileptic-like events. Additionally, distributions in (Fig. 3d-e) are calculated based on the analysis of 1500 realisations of $R(t)$, for each system size and coupling strength indicated. The time-averaged global order parameter $\bar{R}$ in (Fig.4a-d, 7a) is calculated as a long-time average of $5\times 10^{4}$ time steps after a sufficiently large transient of $5\times 10^{4}$ time steps has elapsed, averaged over 50 realisations. The proportion of extreme events $p_{EE}$ in (Fig.7a) is calculated through the analysis of 50 realisations of $A(R(t))$. Simulations in (Fig. 6a-d and 8) consider the time evolution of $1\times 10^{6}$ time steps, and both the distributions and $p_{EE}$ are calculated considering 50 realisations of $A(R(t))$. Finally, the dynamical systems involved in the algorithms for computing the Lyapunov exponents [40, 41, 35] were solved with a fourth-order Runge-Kutta method for $1\times 10^{3}$ time units with an adaptive time step and suitable parameters that ensure convergence.
@@ Appendix B Master Stability Function
To perform a linear stability analysis in order to understand the transition to synchronisation, we study the MSF for the system of Eq. (1) as in [3, 44]. Let $(x_{S},y_{S})^{T}\in\mathcal{S}$, whose orbits obeys:
@@
Let $(\delta x_{i},\delta y_{i})^{T}=(x_{i}(t)-x_{S},y_{i}(t)-y_{S})^{T}$ be small perturbations away from $\mathcal{S}$. Consider now the relation $A_{ij}^{p}=\delta_{ij}k_{i}-\mathcal{L}_{ij}^{p}$, where $\delta_{ij}$ is the Kronecker delta, $k_{i}$ is the degree of the $i$ node, $p$ is the rewiring probability of the Watts-Strogatz model and $A_{ij}^{p},\mathcal{L}_{ij}^{p}$ are the elements of the adjacency and Laplacian matrix of the network $A^{p}$ and $\mathcal{L}^{p}$, respectively. Linearising around $(x_{S},y_{S})^{T}$, and using the property $\sum_{j}\mathcal{L}_{ij}^{p}=0$, the perturbations evolves as
@@
Let now $V^{p}$ be the matrix of eigenvectors of $\mathcal{L}^{p}$ and $\bm{{\xi}}_{i}=(\xi_{i}^{x},\xi_{i}^{y})^{T}$ the projections of the perturbations on the eigenvector basis
@@
which corresponds to the decoupled eigenmodes identified by the eigenvectors of $\mathcal{L}^{p}$ with associated eigenvalues $\gamma_{1}^{p}<\gamma_{2}^{p}<...<\gamma_{i}^{p}<...<\gamma_{N}^{p}$, thus the relation $\mathcal{L}_{kj}^{p}V_{ij}^{p}=\delta_{kj}V_{ij}^{p}\gamma_{i}^{p}$ holds. Using this, Eq. (7) becomes
@@
We compute the MSF of the system which correspond to the maximum Lyapunov exponent $\Lambda_{max}$ of Eq. (12) via the standard procedure developed in [41, 35] for suitable parameters that ensure the convergence of the algorithm.
@@ Appendix C The Wu-Chua conjecture
To quantify and understand in a more detailed manner the stability regions of the synchronisation manifold in the parameter space $N-d$, we test the Wu-Chua conjecture [45], which gives a relation between the second smallest eigenvalue of the Laplacian matrix of a graph and the critical coupling strength needed for synchronisation of two different systems. Let $(s.i)$ and $(s.ii)$ be two linearly coupled systems with $N_{1}$ and $N_{2}$ units with respective laplacian matrices $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ whose smallest eigenvalues are $\gamma_{2}(N_{1})$ and $\gamma_{2}(N_{2})$, then the following relation holds:
@@
where $d^{c}_{N_{1}}$ and $d_{N_{2}}^{c}$ correspond to the critical coupling strength for synchronisation for $(s.i)$ and $(s.ii)$, respectively. We consider the system $(s.i)$ and $(s.ii)$ both to be FitzHugh-Nagumo coupled oscillators whose dynamics obeys Eq. (1). In this case, $(s.i)$ corresponds to the system of $N$ units interacting over a Watts-Strogatz network of the model described in Sec. I with Laplacian matrix $\mathcal{L}_{ij}^{p}$, whose second smallest eigenvalue $\gamma_{2}(N;p,\langle k\rangle)$ is computed numerically, and $(s.ii)$ corresponds to a system of $N_{2}=2$ coupled units whose Laplacian matrix is $\mathcal{L}_{2}=\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$, with second smallest eigenvalue $\gamma_{2}(N_{2})=2$. The synchronisation threshold for system $(s.ii)$ is obtained via the MSF procedure described in Appendix B, and we find it to be $d^{c}_{N_{2}}\approx 0.105\approx 1/10$. The relation (13) becomes
@@
which correspond to the curves shown in Fig. 4b-d and 5b-d, where we have considered $\langle k\rangle=6$ according to the WS network studied.

### v2

@@ Appendix A Numerical simulations
The dynamics of coupled FHN oscillators are solved by integrating Eq. (1) with a fourth-order Runge-Kutta method, considering random initial conditions following a uniform distribution $x\sim U_{x}(-a,a)$, $y\sim U_{y}(-a+a^{3}/3,a+a^{3}/3)$ and a suitable time step that ensures convergence. Specifically, simulations in (Fig. 2) consider the time evolution of $10^{4}$ steps, and the ISI PDF is calculated after a transient of $10^{2}$ time steps. Simulations in (Fig. 3a-e, 4a-d and 7a) consider the time evolution of $10^{5}$ time steps. The computation of $\Omega$ for each point in the $N-d-p$ space in (Fig. 3a-c) is calculated by averaging the analysis of 50 realizations of $R(t)$, excluding the final $10^{3}$ time steps, in order to consider only ESR events. Additionally, distributions in (Fig. 3d-e) are calculated based on the analysis of 1500 realizations of $R(t)$, for each system size and coupling strength indicated. The time-averaged global order parameter $\bar{R}$ in (Fig.4a-d, 7a) is calculated as a long-time average of $5\times 10^{4}$ time steps after a sufficiently large transient of $5\times 10^{4}$ time steps, averaged over 50 realizations. The proportion of EEs $p_{EE}$ in (Fig.7a) is calculated through the analysis of 50 realizations of $A(R(t))$. Simulations in (Fig. 6a-d and 8) consider the time evolution of $10^{6}$ time steps, and both the distributions and $p_{EE}$ are calculated considering 50 realizations of $A(R(t))$. Finally, the dynamical systems involved in the algorithms for computing the Lyapunov exponents [52, 53, 47] were solved with a fourth-order Runge-Kutta method for $10^{3}$ time units with an adaptive time step and suitable parameters that ensure convergence.
@@ Appendix B Master Stability Function
To perform the linear stability analysis for understanding the transition to synchronization, we study the MSF of the system of Eq. (1) as in [3, 56]. Let $(x_{S},y_{S})^{T}\in\mathcal{S}$, whose orbits obey:
@@
Let $(\delta x_{i},\delta y_{i})^{T}=(x_{i}(t)-x_{S},y_{i}(t)-y_{S})^{T}$ be small perturbations away from $\mathcal{S}$. Consider now the relation $A_{ij}^{p}=\delta_{ij}k_{i}-\mathcal{L}_{ij}^{p}$, where $\delta_{ij}$ is the Kronecker delta, $k_{i}$ is the degree of the $i$ node, $p$ is the rewiring probability of the Watts-Strogatz model and $A_{ij}^{p},\mathcal{L}_{ij}^{p}$ are the elements of the adjacency and Laplacian matrix of the network $A^{p}$ and $\mathcal{L}^{p}$, respectively. Linearizing around $(x_{S},y_{S})^{T}$, and using the property $\sum_{j}\mathcal{L}_{ij}^{p}=0$, the perturbations evolves as
@@
Now let $V^{p}$ be the matrix of eigenvectors of $\mathcal{L}^{p}$ and $\bm{{\xi}}_{i}=(\xi_{i}^{x},\xi_{i}^{y})^{T}$ the projections of the perturbations on the eigenvector basis
@@
which corresponds to the decoupled eigenmodes identified by the eigenvectors of $\mathcal{L}^{p}$ with associated eigenvalues $\gamma_{1}^{p}<\gamma_{2}^{p}<...<\gamma_{i}^{p}<...<\gamma_{N}^{p}$, thus, the relation $\mathcal{L}_{kj}^{p}V_{ij}^{p}=\delta_{kj}V_{ij}^{p}\gamma_{i}^{p}$ holds. Using this, Eq. (7) becomes
@@
We compute the MSF of the system, which corresponds to the maximum Lyapunov exponent $\Lambda_{max}$ of Eq. (12) via the standard procedure developed in [53, 47] for suitable parameters that ensure the convergence of the algorithm.
@@ Appendix C The Wu-Chua conjecture
To quantify and understand in a more detailed manner the stability regions of the synchronization manifold in the $N-d$ parameter space, we test the Wu-Chua conjecture [57], which gives a relation between the second smallest eigenvalue of the Laplacian matrix of a graph and the critical coupling strength needed for synchronization of two different systems. Let $(s.i)$ and $(s.ii)$ be two linearly coupled systems with $N_{1}$ and $N_{2}$ units with respective Laplacian matrices $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ whose smallest eigenvalues are $\gamma_{2}(N_{1})$ and $\gamma_{2}(N_{2})$, then the following relation holds:
@@
where $d^{c}_{N_{1}}$ and $d_{N_{2}}^{c}$ correspond to the critical coupling strength for synchronization for $(s.i)$ and $(s.ii)$, respectively. We consider the system $(s.i)$ and $(s.ii)$ both to be FitzHugh-Nagumo coupled oscillators whose dynamics obeys Eq. (1). In this case, $(s.i)$ corresponds to the system of $N$ units interacting over a Watts-Strogatz network of the model described in Sec. I with Laplacian matrix $\mathcal{L}_{ij}^{p}$, whose second smallest eigenvalue $\gamma_{2}(N;p,\langle k\rangle)$ is computed numerically, and $(s.ii)$ corresponds to a system of $N_{2}=2$ coupled units whose Laplacian matrix is $\mathcal{L}_{2}=\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$, with second smallest eigenvalue $\gamma_{2}(N_{2})=2$. The synchronization threshold for system $(s.ii)$ is obtained via the MSF procedure described in Appendix B, and we find it to be $d^{c}_{N_{2}}\approx 0.105\approx 1/10$. The relation (13) becomes
@@
which correspond to the curves shown in Fig. 4b-d and 5b-d, where we have considered $\langle k\rangle=6$ according to the model studied.